OpenStudy (anonymous):

five dice are rolled.find the probability that three dice show the composite number

OpenStudy (rane):

do u know what is composite no's ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help

OpenStudy (anonymous):

composite no are those no. with are not prime..

OpenStudy (anonymous):

4,6,8,10,12,9,etc

OpenStudy (anonymous):

2^6/6^6

OpenStudy (anonymous):

yep right.. and i am not sure about the answer

OpenStudy (anonymous):

64/7776

OpenStudy (anonymous):

nope in the numerator something wrong

OpenStudy (anonymous):

It's a binomial distribution. \(p=1/3\). since 2 of 6 sides are composite. Also \(1-p=2/3\),\(k=3\), and \(n=5\) \[ \Pr(k=3) = \binom{5}{2}(1/3)^3(1-1/3)^{5-3} \] That gives me \(40/243\approx 16.5\%\) http://www.wolframalpha.com/input/?i=%285+choose+2%29%281%2F3%29%5E3%281-1%2F3%29%5E%7B5-3%7D

OpenStudy (anonymous):

The binomial distribution is as follows. You have some event which succeeds with probability \(p\). In this case the event is a dice roll. You need repeat the trial \(n\) times and \(X\) is the random variable representing the number of successes. \[ \Pr(X=k)=\binom{n}{k}p^k(1-p)^{n-k} \]

OpenStudy (anonymous):

In this case we wanted \(X=k=3\) successes out of the \(n=5\) trials. The success probability was \(p=2/6=1/3\) because there are two composite numbers \(4\) and \(6\) out of all number on a die: \(1,2,3,4,5,6\).

OpenStudy (anonymous):

how 2 of 6 are composite.plse explain @wio

OpenStudy (anonymous):

\(4=2\times 2\) and \(6=3\times 2\)

OpenStudy (anonymous):

\(2\), \(3\), and \(5\) are prime. \(1\) is not composite.

OpenStudy (anonymous):

k understood wio.i have a confusion always btw a permutations ,probability etc can u explain .how to know permutations and probability. and binomial distribution

OpenStudy (anonymous):

I already explain binomial distribution.

OpenStudy (anonymous):

k .then permutation and probability

OpenStudy (anonymous):

I'm tired.

OpenStudy (anonymous):

its k .tell me the website to learn these topics clearly

OpenStudy (anonymous):

thank u very much.

OpenStudy (anonymous):

do u know cat exam